### v2

# Curious crossing-critical edges – variations on an example of S̆irán̆

Abstract.
Motivated by Kuratowski’s theorem, a Kuratowski subgraph of a graph is a subgraph that is a subdivided $K_{5}$ or a subdivided $K_{3,3}$. An edge is crossing-critical if the crossing number decreases after removing the edge. In this note, we present the following examples: a graph with an edge that is crossed in every optimal drawing of the graph, but the edge is not in any Kuratowski subgraph of the graph; a graph with an edge that is in every Kuratowski subgraph but is not crossed in any optimal drawing of the graph; and a graph with a crossing-critical edge that is not present in any Kuratowski subgraph and is not crossed in any optimal drawing of the graph. Fáry’s theorem implies that the Kuratowski subgraphs are the only obstructions to a graph having a crossing-free drawing with all edges drawn as straight lines. The three example graphs given also hold if we restrict drawings to only have straight line edges, and thus also apply to the rectilinear crossing number.
Key words and phrases:
graph drawing, crossing number, crossing-critical, Kuratowski subgraph
2020 Mathematics Subject Classification:
Primary 05C10
1. Introduction
In this note, a graph is a finite loopless multigraph, i.e., we allow parallel edges but no loops. Edges are parallel if they have the same endvertices. Given vertices $x,y$ of the graph $G$, we will use $m_{xy,G}$ for the number of $x$-$y$ edges in $G$; we call $m_{xy,G}$ the multiplicity of the $x$-$y$ edge in $G$. A simple graph is a graph $G$ without parallel edges, i.e., a graph where $m_{xy,G}=1$ for each adjacent vertex pair $x,y$. Given a multigraph $G$, the underlying simple graph is a simple graph $H$ on the same vertex set, such that $H$ has a single edge between vertices $x,y$ precisely when $x,y$ are adjacent in $G$. If $G$ is a graph, a representation of $G$ is the underlying simple graph $H$ together with a weighting of the edges, where the edge weight of the $x$-$y$ edge of $H$ is $m_{xy,G}$.
A subdivision of a graph $H$ is a graph $G$ where some of the edges of $H$ are replaced by paths whose internal vertices are not in $H$ and not shared with any other path. The vertices of $H$ are called the main vertices of the subdivision, and the internal vertices of the paths are called the subdividing vertices.
A drawing $\mathcal{D}$ of a graph $G$ is an embedding of the graph in the plane such that the vertices are assigned to points of the plane injectively, the edges are mapped to Jordan arcs connecting to the endvertices of the edge, and if $v$ is not an endvertex of the edge $e$ then the image $\mathcal{D}(v)$ of $v$ is disjoint from the image $\mathcal{D}(e)$ of $e$.
A graph is planar if it can be drawn in the plane without crossing edges, i.e., such that the images of two edges only have their common endvertices in common; in other words edges do not cross in the drawing. The celebrated Kuratowski theorem [3] states that a graph if planar precisely when it does not contain a subdivision of the complete graph $K_{5}$ or the complete bipartite graph $K_{3,3}$. $H$ is called a Kuratowski subgraph of the graph $G$ if $H$ is a subgraph of $G$ and $H$ is a subdivision of either $K_{3,3}$ or $K_{5}$. Note that Kuratowski subgraphs are simple graphs by definition. In a sense, Kuratowski theorem states that Kuratowski subgraphs are forcing crossings to appear in any drawing of the graph. Fáry’s theorem (proved independently in [2, 6]) states that every planar graph can be drawn without crossings even if we restrict to drawings using straight line edges (the rectilinear drawings). Thus, the Kuratowski subgraphs are also the only obstacles to having planar straight line drawings.
The crossing number $\operatorname{cr}(\mathcal{D})$ of a drawing $\mathcal{D}$ of $G$ is $\sum\limits_{e,f\in E(G)}|\mathcal{D}(e)^{0}\cap\mathcal{D}(f)^{0}|$, where for an edge $e$, $\mathcal{D}(e)^{0}$ is the interior of $\mathcal{D}(e)$ , and the summation goes for unordered pairs of distinct edges. The crossing number $\operatorname{cr}(G)$ of a graph $G$ is the minimum crossing number over all its drawings. A drawing $\mathcal{D}$ of $G$ is called optimal if $\operatorname{cr}(\mathcal{D})=\operatorname{cr}(G)$. An edge $e\in E(G)$ is called crossing-critical if $\operatorname{cr}(G-e)<\operatorname{cr}(G)$. The rectilinear crossing number $\overline{\operatorname{cr}}(G)$ of the graph $G$ is the minimum crossing number of its rectilinear drawings. Clearly $\operatorname{cr}(G)\leq\overline{\operatorname{cr}}(G)$, with equality precisely when $G$ has an optimal drawing that is rectilinear. An edge $e\in E(G)$ called crossing-critical with respect to the rectilinear crossing number if $\overline{\operatorname{cr}}(G-e)<\overline{\operatorname{cr}}(G)$.
If $h$ is an edge that is crossed in some optimal drawing $\mathcal{D}$ of $G$, then $h$ is clearly crossing-critical, as the drawing of GhG-h induced by $\mathcal{D}$ has fewer crossings than $\operatorname{cr}(G)$.
If $G$ is a non-planar graph and $h$ is an edge that belongs to every Kuratowski subgraph of $G$, then $h$ is crossing-critical, as by Kuratowski’s theorem $\operatorname{cr}(G-h)=0$, and by Fáry’s theorem this edge is also crossing critical with respect to the rectilinear crossing number.
The purpose of this note is the following:
Problem 1.
Find the answer to the following questions:
  1. (1)
    If an edge $h$ of a graph $G$ is crossed in every optimal drawing of $G$, must $h$ be in some Kuratowski subgraph?
  2. (2)
    If $G$ is a graph with an edge $h$ such that $h$ is in every Kuratowski subgraph, must $G$ have an optimal drawing in which $h$ is crossed?
  3. (3)
    If the edge $h$ of a graph $G$ is not crossed in any optimal drawing of $G$ and is not present in any Kuratowski subgraph of $G$, is it true that edge $h$ cannot be crossing-critical?
[…]
Figure 1. S̆irán̆’s example [5]: $\operatorname{cr}(G^{*}-e)<\operatorname{cr}(G^{*})=6$, $e$ is not in any Kuratowski subgraph of $G^{*}$, but is crossed in the optimal drawing shown.
In 1983, S̆irán̆ [5] gave an example of a graph $G^{*}$ that has a crossing-critical edge $e$ such that $e$ is not an edge of any of the Kuratowski subgraphs of $G^{*}$. Figure 1 shows an optimal drawing of this graph $G^{*}$, in which $e$ is crossed, its underlying simple graph $G_{0}$ and a representation of $G^{*}$. The graph $G^{*}$ is not a negative answer to our first question, as $G^{*}$ does have optimal drawings in which $e$ is not crossed. We modify S̆irán̆’s example by changing the edge-multiplicities of $G^{*}$ judiciously to create graphs $G_{1},G_{2},G_{3}$ (see Figure 2) such that each have the same underlying simple graph $G_{0}$ that give the following (perhaps surprising) answers to our questions:
Claim 1.
The answer to all questions in Problem 1 is in the negative (see Figure 2 for the examples), i.e.,
  1. (1)
    A graph can have an edge that is crossed in every optimal drawing, but is not in any Kuratowski subgraph of the graph. (An example is the graph $G_{1}$ and the unique $s$-$t$ edge $e$.)
  2. (2)
    A non-planar graph can have an edge that is in every Kuratowski subgraph but is not crossed in any optimal drawing of the graph. (An example is the graph $G_{2}$ and the unique $v_{1}$-$v_{3}$ edge $f$.)
  3. (3)
    A graph can have an edge that is crossing-critical, but this edge is not present in any Kuratowski subgraph and is not crossed in any optimal drawing of the graph. (An example is the graph $G_{3}$ and the unique $s$-$t$ edge.)
Furthermore, all three negative answers can be provided with simple graphs. Moreover, the these simple graphs provide the corresponding examples for the rectilinear crossing number as well.
TWe note that the “furthermore” part of the previous claim is obvious: in each of the $G_{i}$, subdivide all but one of the parallel copies of each $x$-$y$ edge to obtain a simple graph $G_{i}^{*}$. The edge $e$ does not have parallel copies, so it is never subdivided, and the edge $f$ does not have parallel copies in $G_{2}$. Hence these simple graphs also have the required properties. The “moreover” part follows from the fact that we have straight line drawings of the underlying simple graphs $G_{i}$ (Figure 3) used in the proof, and the simple graphs $G_{i}^{*}$ have corresponding straight line drawing: draw the not subdivided copy of a multi-edge as a straight line, and by placeing the subdivideding copies of this edges of the multi-edges close to the original edge.
[…]
Figure 2. The graphs $G_{1}$ and $G_{2}$ and a representation of $G_{3}$. $\ell_{1}=8$, $\ell_{2}=2$ and $\ell_{3}=22$.
Note that the graphs $G_{1},G_{2},G_{3}$ all have $G_{0}$ as the underlying simple graph, in which we specifically some of the edges (e.g. $e$). We will use the following convention: Let $p\in\{e,f,g,h\}$ and $m_{p,G_{i}}=k$. For $k\geq 1$, we denote the parallel edges represented by $p$ as $p_{1},p_{2},\ldots,p_{k}$, but if $k=1$, we just call the edge $p$. Note that for each $i\in\{1,2,3\}$ there is an integer $\ell_{i}$ such that for any edge $p$ of $G_{0}$, if $p\notin\{e,f,g,h\}$ then $m_{p,G_{i}}=\ell_{i}$, and for $p\in\{e,f,g,h\}$, we have $m_{p,G_{i}}<\ell_{i}$. Note that $\ell_{1}=8$, $\ell_{2}=2$ and $\ell_{3}=22$ (see Figure 2). We will use these $\ell_{i}$ quantities later.
We will show that the graphs $G_{1},G_{2},G_{3}$ and the edges $e,f$ satisfy the properties claimed in Claim 1 by first finding the edges of Kuratowski subgraphs of the graphs in Claim 2, and then (after some preliminary work) verifying the rest of Claim 1 in Claim 5.
2. Proof of our claims
First we determine which edges belong to Kuratowski subgraphs.
Claim 2.
For $i\in\{0,1,2,3\}$, let $F_{i}$ denote the set of edges of $G_{i}$ that are present in some Kuratowski subgraph of $G_{i}$. Then $F_{i}=E(G_{i})-\{e\}$, so $e$ is not in any Kuratowski subgraph of $G_{i}$. Moreover, if $i\neq 3$, then $f$ is present in every Kuratowski subgraph of $G_{i}$.
Proof.
Assume first that $i\in\{0,1,2\}$. The drawings in Figure 2 illustrate that $G_{i}-f$ is planar, therefore $f$ is present in every Kuratowski subgraph of $G_{i}$.
Now let $i\in\{1,2,3\}$. If a Kuratowski subgraph $H$ of $G_{i}$ contains an edge $p$, and $p^{\prime}$ is parallel to $p$, then $(H-p)+p^{\prime}$ is also a Kuratowski subgraph of $G_{i}$. Thus, it is enough to show that $F_{0}=E(G_{0})-\{e\}$.
As any subdivision of $K_{5}$ has $5$ vertices of degree $4$, and $G_{0}$ has only $3$ vertices of degree at least $4$, any Kuratowski subgraph of $G_{0}$ is a subdivided $K_{3,3}$ with $6$ vertices of degree $3$, and at most $1$ subdividing vertex of degree $2$.
We define the main vertex classes of the $K_{3,3}$ Kuratowski subgraph as the two disjoint 3-element sets of main vertices that correspond to the partite sets of the $K_{3,3}$ after suppressing the subdividing vertices.
Let $H$ be a Kuratowski subgraph of $G_{0}$. This $H$ is a subdivided $K_{3,3}$. First, suppose $H$ has no subdividing vertices. Then there is a vertex $y\in V(G_{0})$ such that $H$ is a spanning subgraph of $G_{0}-y$. As $H$ is $3$-regular, $y$ is not adjacent to any degree $3$ vertex in $G_{0}$. Therefore $y\in\{v_{2},v_{4}\}$. As for $j\in\{2,4\}$ the graph $G_{0}-v_{j}$ is $3$-regular, so we now have $H=G-v_{j}$. However, $G-v_{j}$ contains the $3$-cycle $uv_{1}v_{3}$ and therefore it is not $K_{3,3}$. Thus, $H$ must have a unique subdividing vertex $w$.
If $w\notin\{v_{1},v_{3}\}$ then $v_{1}$ and $v_{3}$ are main (and therefore degree $3$) vertices of $H$. Therefore all of the $3$ edges coming out of each of them must be in $H$. Thus, $v_{1}$ and $v_{3}$ are connected by an edge in $H$, consequently they are not in the same main vertex class. Moreover, as $u$ is a neighbor of both $v_{1}$ and $v_{3}$, if $u$ is a subdividing vertex, then the $v_{1}uv_{3}$ path is a subdivided $v_{1}v_{3}$ edge. However, the $v_{1}v_{3}$ edge of $G_{0}$ is already in $H$, which gives that $u$ is a main vertex of $H$. This gives a triangle on the main vertices $u,v_{1},v_{3}$, a contradiction. Therefore we must have $w=v_{i}$ for some $i\in\{1,3\}$, and the main vertices of $H$ are $s,t,u,v_{2},v_{4},v_{4-i}$.
If $w=v_{1}$, then, as the neighbors of $v_{3}$ are $t,u$ and $v_{1}$, one of the main vertex classes of $H$ must be $\{t,u,z\}$ where $z$ is a neighbor of $v_{1}$ different from $v_{3},t,u$. Therefore the main vertex classes of $H$ are $\{t,u,s\}$ and $\{v_{2},v_{4},v_{3}\}$, and $H=G_{0}-\{e,r\}$. If $w=v_{3}$, then, as the neighbors of $v_{1}$ are $s,u$ and $v_{3}$, one of the main vertex classes of $H$ must be $\{s,u,z\}$, where $z$ is a neighbor of $v_{3}$ different from $v_{1},s,u$. Therefore the main vertex classes of $H$ are $\{s,u,t\}$ and $\{v_{2},v_{4},v_{1}\}$, and $H=G_{0}-\{e,q\}$.
This gives that $F_{0}=E(G_{0})-\{e\}$, which finishes our proof. ∎
Note that it is easy to determine from the proof above that the number of Kuratowski subgraphs in $G_{1}$ is $2\cdot 8^{7}\cdot 3^{2}=37,748,736$, in $G_{2}$ is $2^{8}=256$ and in $G_{3}$ is $2^{2}\cdot 22^{7}\cdot 3\cdot 7=209,526,062,592$.
Next we consider the drawings. It is well-known [4] that any optimal drawing of a graph must be good, i.e., it must be such that any two arcs representing edges in the drawing do not touch each other, they cross at most once, and arcs representing edges that share an endvertex do not cross.
If $H$ is an edge-weighted simple graph with edge-weights $w:E(G)\rightarrow\mathbb{R}^{+}$, the weighted crossing number $\operatorname{cr}_{w}(\mathcal{D})$ of a good drawing $\mathcal{D}$ is
$\operatorname{cr}_{w}(\mathcal{D})=\sum_{e,f\text{ cross in}\mathcal{D}}w(e)w(f).$
and $\operatorname{cr}_{w}(H)$ is the minimum of the weighted crossing numbers of good drawings of $H$. (Note that we need not restrict ourselves to good drawings in the definition of weighted crossing number, we do so only to eliminate unnecessary complications in the definition).
We will make use of the following easy claim:
Claim 3.
Let $G$ be any graph, $\mathcal{D}$ be an optimal drawing of $G$, and let $H$ be the underlying simple graph of $G$, and let $w$ be the edge-weighting of the representation of $G$. If $e\in E(G)$ is an $x$-$y$ edge, let $e_{H}$ denote the unique $x$-$y$ edge in $H$. The following statements are true:
  1. (i)
    If $e,e^{\prime}$ are parallel edges of $G$, then $e,e^{\prime}$ have the same number of crossings in $\mathcal{D}$.
  2. (ii)
    For any edge $e$ of $G$ there is an optimal drawing $\mathcal{D}^{\prime}_{e}$ of $G$ such that $\mathcal{D}$ and $\mathcal{D}^{\prime}_{e}$ only differ on edges parallel to $e$ but differing from $e$, and if an edge $e^{\prime}$ is parallel with $e$, then $e^{\prime}$ crosses the same edges in $\mathcal{D}^{\prime}_{e}$ as $e$ does in $\mathcal{D}$.
  3. (iii)
    For any edge $e$ of $G$ there is an optimal drawing $\mathcal{D}_{e}$ of $G$ such that any two parallel edges cross exactly the same edges in $\mathcal{D}_{e}$, and if $e$ crosses an edge in $\mathcal{D}$, then it crosses that edge in $\mathcal{D}_{e}$ as well.
  4. (iv)
    Let $e$ be an $x$-$y$ edge and $f$ be a $z$-$w$ edge of $G$. If $e,f$ cross in $\mathcal{D}$, then $m_{xy}m_{zw}\leq\operatorname{cr}(G)$.
  5. (v)
    For any edge $e$ of $G$ there is a good drawing $\overline{\mathcal{D}}_{e}$ of $H$ such that $\operatorname{cr}(G)=\operatorname{cr}_{w}\left(\overline{\mathcal{D}}_{e}\right)$ and if for some edge $g$ of $G$, $e$ and $g$ cross in $\mathcal{D}$ then $e_{H},g_{H}$ cross in $\overline{\mathcal{D}}_{e}$.
  6. (vi)
    $\operatorname{cr}(G)=\operatorname{cr}_{w}(H)$.
Proof.
Assume that, contrary to (i), there are parallel edges $e,e^{\prime}$ such that they have different number of crossings in $\mathcal{D}$; without loss of generality $e$ has fewer crossings than $e^{\prime}$. Obtain a new drawing $\mathcal{D}^{\prime}$ by keeping all but the edge-curve of $e^{\prime}$, and redraw $e^{\prime}$ close to $e$ so that the curve of $e^{\prime}$ intersect exactly those edges that $e$ intersects. Then $\operatorname{cr}(\mathcal{D}^{\prime})<\operatorname{cr}(\mathcal{D})$, a contradiction. (i) follows.
Now fix the edge $e$ and create the drawing $\mathcal{D}^{\prime}_{e}$ by changing only the drawing of the edges $e^{\prime}$ in $\mathcal{D}$ that are parallel with $e$ by redrawing these edges close to $e$ such that they intersect the same edges as $e$ did. This shows (ii) holds,
To show (iii), we apply the procedure in (ii) repeatedly as follows: Let $h_{1}=e_{H},h_{2},\ldots,h_{m}$ be an enumeration of the edges of $H$, and chose the edges $g_{1}=e,g_{2},\ldots,g_{m}$ such that $h_{i}=g_{i,H}$ for all $i\in[m]$ and $g_{i}$ is chosen such that if $e$ intersects any copies of $h_{i}$ in $\mathcal{D}$, then $g_{i}$ is one of these intersecting edges. Define the sequence of drawings $\mathcal{D}_{0}=\mathcal{D},\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{m}$ such that for all $i\in[m]$, $\mathcal{D}_{i}=(\mathcal{D}_{i-1})^{\prime}_{g_{i}}$, i.e., $\mathcal{D}_{i}$ is obtained from $\mathcal{D}_{i-1}$ by redrawing the edges parallel to $g_{i}$. Easy induction shows that for each $i\in[m]$, $\mathcal{D}_{i}$ satisfies the following: Let $f$ be an edge of $G$, and let $j$ be the index such that $f_{H}=h_{j}$. If $j>i$ then $f$ and $e$ cross in $\mathcal{D}_{i}$ precisely when $f$ and $e$ cross in $\mathcal{D}$. If $j\leq i$ then $f,f^{\prime}$ cross the same edges in $\mathcal{D}_{i}$ for all edges $f^{\prime}$ that are parallel to $f$ and if $f,e$ cross in $\mathcal{D}$, then they cross in $\mathcal{D}_{i}$. Selecting for $\mathcal{D}_{e}$ the drawing $\mathcal{D}_{m}$, the claims of (iii) are satisfied.
Part (iv) follows from (iii).
Now for an edge $e$ of $G$ consider the drawing $\mathcal{D}_{e}$ in (iii), for each edge of $H$ select a representative edge among their parallel copies, and set $\overline{\mathcal{D}}_{e}$ be the subdrawing of $\mathcal{D}_{e}$ induced by these edges. This drawing clearly satisfies the claims of  (v).
Finally, to prove (vi)(v) implies $\operatorname{cr}(G)\geq\operatorname{cr}_{w}(H)$. Let $\overline{\mathcal{D}}$ be a $\operatorname{cr}_{w}$-optimal drawing of $H$, and create a drawing $\mathcal{D}^{\prime}$ of $G$ by drawing all parallel edges close to the drawing of their copy in $\overline{\mathcal{D}}$ such that they do not cross each other, cross only copies of edges that their copy crosses, and they cross every edge at most once. Then $\operatorname{cr}(G)\leq\operatorname{cr}(\mathcal{D}^{\prime})=\operatorname{cr}_{w}(\overline{\mathcal{D}})\leq\operatorname{cr}_{w}(H)$, finishing the proof. ∎
We detour a bit by defining when two drawings of a graph are equivalent (see standard graph theory books, e.g., [1]). Using $\mathcal{D}(G)$ for the set of points that are in the drawing, we can define a face of the drawing as a maximal connected component of $\mathbb{R}^{2}-\mathcal{D}(G)$ (note that this does not assume that the drawing is planar).
Let $S^{2}$ be the two-dimensional sphere, with north pole $(0,0,1)$, and $\pi:S^{2}-\{(0,0,1)\}\rightarrow\mathbb{R}^{2}$ is the stereographic projection of the sphere to the plane. Let $G$ be a graph and $\mathcal{D}_{1},\mathcal{D}_{2}$ be two drawings of $G$. Drawings $\mathcal{D}_{1},\mathcal{D}_{2}$ are called equivalent if there is a homeomorphism $\phi:S^{2}\rightarrow S^{2}$ such that $\pi\circ\phi\circ\pi^{-1}$ maps $\mathcal{D}_{1}$ to $\mathcal{D}_{2}$. Note that the use of $\pi,\pi^{-1}$ allows us to switch the outer face of the drawing without changing any salient features of the drawing.
Recall that for $k\in\mathbb{N}$ a simple graph is $k$-connected if it has more than $k$ vertices and the removal of fewer than $k$ vertices does not disconnect the graph. Whitney [7] showed that any two crossing-free drawings of a $3$-connected simple planar graph are equivalent.
Claim 3 allows us to reduce our questions about optimal drawing of a graph into questions about drawings of the underlying simple graph. So, we will first consider drawings of the simple graph $G_{0}$.
Claim 4.
Assume that $\mathcal{D}$ is a drawing of $G_{0}$ such that the induced subdrawing of $G_{0}-f$ is planar, and the set of edges that $f$ crosses in $\mathcal{D}$ is a subset of $\{e,g,h\}$. Then $f$ crosses each of $e$,$g$ and $h$ in $\mathcal{D}$.
Proof.
Recall the drawing $\mathcal{D}_{0}$ of $G_{0}$ (Figure 1). $G_{0}-f$ has two vertices of degree $2$, $v_{1}$ and $v_{3}$. Let $H_{0}$ be the simple graph obtained from $G_{0}-f$ by suppressing $v_{1}$ and $v_{3}$ as subdividing vertices. Then $V(H_{0})=\{t,u,s,v_{2},v_{4}\}$ and any two vertices are adjacent in $H_{0}$ except $v_{2}$ and $v_{4}$ (in other words, $H$ is a $K_{5}$ minus an edge). The simple graph $H_{0}$ is 3-connected, so Whitney’s theorem applies, and the subdrawings $\mathcal{D}_{0}|_{H_{0}},\mathcal{D}|_{H_{0}}$ of $H_{0}$ induced by the drawings $\mathcal{D}_{0},\mathcal{D}$ are equivalent. This means that there is a drawing $\mathcal{D}^{\prime}$ of $G_{0}$ that is equivalent to $\mathcal{D}$ such that $\mathcal{D}_{0}|_{G_{0}-f}=\mathcal{D}^{\prime}|_{G_{0}-f}$, and to prove our claim it is enough to show that $f$ crosses all of $e,g,h$ in $\mathcal{D}^{\prime}$. The edges $E(G_{0})-\{e,f,g,h\}$ form the closed walk $v_{3}tv_{2}uv_{1}sv_{4}uv_{3}$. The subdrawings of this closed walk in $\mathcal{D}^{\prime}$ (and $\mathcal{D}_{0}$) divide the plane into the same $3$ regions, one of which contains the edges of $e,g,h$ (Figure 1, middle picture). As $G$ is non-planar, $f$ must cross some edges in $\mathcal{D}^{\prime}$ (and $\mathcal{D}_{0}$). Since $f$ can only cross the edges of $e,g,h$, edge $f$ must also lie in the same region as $e,g,h$, and must cross one of them. However, if it crosses one of them, it crosses all of them. ∎
[…]
Figure 3. The rectilinear drawing $\mathcal{D}_{4}$ of $G_{0}$, and a rectilinear drawing equivalent with $\mathcal{D}_{0}$.
Now we are ready to prove our final claim:
Claim 5.
The following statements are true:
  1. (i)
    For all $i\in\{1,2,3\}$, we have $\operatorname{cr}(G_{i})\leq\ell_{i}-1$.
  2. (ii)
    $\operatorname{cr}(G_{1})=7=\ell_{1}-1$, and $e$ is crossed in every optimal drawing of $G_{1}$.
  3. (iii)
    $\operatorname{cr}(G_{2})=1=\ell_{2}-1$, and $f$ is not crossed in any optimal drawing of $G_{2}$.
  4. (iv)
    $\operatorname{cr}(G_{3})=21=\ell_{3}-1$, and $e$ is not crossed in any optimal drawing of $G_{3}$, but $\operatorname{cr}(G-e)<\operatorname{cr}(G)$.
Proof.
For $i\in\{1,2,3\}$, let $w_{i}$ denote the edge-weights of $G_{0}$ in the representation of $G_{i}$
The drawing $\mathcal{D}_{1}$ on Figure 2 yields $\operatorname{cr}(G_{1})\leq 7=\ell_{1}-1$. The drawing $\mathcal{D}_{4}$ on Figure 3 together with Claim 3 (vi) gives that for $i\in\{2,3\}$, $\operatorname{cr}(\mathcal{D}_{i})\leq\operatorname{cr}_{w_{i}}(\mathcal{D}_{4})=w_{i}(h)w_{i}(g)=\ell_{i}-1$. Therefore part (i) of this claim is true.
Let $i\in\{1,2,3\}$, and let $\mathcal{D}^{*}_{i}$ be any optimal drawing of $G_{i}$.
Since $\operatorname{cr}(G_{i})\leq\ell_{i}-1$, Claim 3 (iv) gives that only copies of $e,f,g$ or $h$ may cross in $\mathcal{D}^{*}_{i}$. Since $\mathcal{D}^{*}_{i}$ is a good drawing, a crossing pair of edges are between a copy of $f$ and a copy of one of $e,g,h$, or between a copy of $h$ and a copy of $g$ in $G_{i}$.
Since $w_{1}(g)w_{1}(h)=3\cdot 3>7\geq\operatorname{cr}(G_{1})$, by Claim 3 (iv) no optimal drawing of $G_{1}$ has a crossing between a copy of $g$ and a copy of $h$. This yields that in every optimal drawing of $G_{1}$, all crossings involve $f$ and a copy of one of $e,g,h$. Take an arbitrary copy of each of the parallel edges, and consider the subdrawing of this graph induced by $\mathcal{D}^{*}_{1}$. Claim 4 gives that $f$ crosses all copies of $e,g,h$ in this subdrawing. Therefore $f$ crosses all copies of $e,g,h$ in $\mathcal{D}^{*}_{i}$, $\operatorname{cr}(\mathcal{D}^{*}_{1})=w_{1}(h)+w_{1}(g)+w_{1}(e)=7$, and $e$ is crossed in any optimal drawing of $G_{1}$. This finishes the proof of part (ii).
Clearly $\operatorname{cr}(G_{2})=1$, as $G_{2}$ has at least one Kuratowski subgraph and $\operatorname{cr}(G_{2})\leq 1$. Thus if $f$ is crossed $\mathcal{D}^{*}_{2}$, all crossings of $\mathcal{D}^{*}_{2}$ are on $f$. But by Claim 4 this gives that $f$ crosses all copies of $e,g,h$, which contradicts $\operatorname{cr}(G_{2})=1$. This finishes the proof of (iii).
We move to the proof of case (iv). If no copy of $g$ crosses a copy of $h$ in $\mathcal{D}^{*}_{3}$, then all crossings involve a copy of $f$ and a copy of one of $e,h,g$. Using Claim 4 we get, as before, that all copies of $f$ cross all copies of $e,h,$ and $g$ in $\mathcal{D}^{*}_{3}$, giving $\operatorname{cr}(\mathcal{D}^{*}_{3})=w_{3}(f)(w_{3}(e)+w_{3}(h)+w_{3}(g))=22>\operatorname{cr}(G_{2})$, a contradiction. Therefore we have that some copy of $g$ crosses some copy of $h$ in every optimal drawing of $G_{3}$. Together with Claim 3 (iv) this yields $21\geq\operatorname{cr}(G_{3})\geq w_{3}(h)w_{3}(g)=21$.
Consider the drawing $\mathcal{D}_{0}$ of $G_{0}$ (see Figure 1). Using the subdrawing of $G_{0}-e$ induced by this drawing and Claim 3 (vi) gives that $\operatorname{cr}(G_{3}-e)\leq\operatorname{cr}_{w_{3}}(G_{0}-e)\leq w_{3}(f)(w_{3}(g)+w_{3}(h))=20<\operatorname{cr}(G_{3})$.
To finish the proof of (iv), we only need to show that $e$ is not crossed in any optimal drawing of $G_{3}$. Note that we already know that if $e$ is crossed in an optimal drawing, it crosses $f_{1}$ or $f_{2}$.
Set $N=\{\{h_{i},g_{j}\}:h_{i}\text{ does not cross }g_{j}\text{ in }\mathcal{D}^{*}_{3}\}$.
If $N\neq\emptyset$, let $\{h_{i},g_{j}\}\in N$. Let $H$ be the simple graph containing $f_{1},h_{i},g_{j}$, which contains exactly one copy of each parallel edges of $G_{3}$. Clearly, $H$ is isomorphic of $G_{0}$. The subdrawing $\overline{\mathcal{D}^{*}_{3}}$ of $H$ induced by $\mathcal{D}^{*}_{3}$ only has crossings between $f_{1}$ and $h_{i},g_{j},e$. Therefore, Claim 4 gives that $f_{1}$ crosses $h_{i},g_{j},e$ in $\mathcal{D}^{*}_{3}$ . Using Claim 3 (iii) there is an optimal drawing $\mathcal{D}^{*}_{3,f_{1}}$ such that both $f_{1},f_{2}$ cross all copies of $h,g$ in $G_{3}$. However, this means $\operatorname{cr}(\mathcal{D}^{*}_{3})=\operatorname{cr}(\mathcal{D}^{*}_{3,f_{1}})\geq 22$, a contradiction. Thus we have $N=\emptyset$.
If $N=\emptyset$, all copies of $h$ cross all copies of $g$ in $\mathcal{D}^{\prime}_{3}$, and the copies of $h$ and $g$ produce all $21$ crossings of the drawing, consequently other edges (in particular $e$) are not involved in a crossing. ∎
It is clear that Claim 2 and Claim 5 implies Claim 1.